# A new framework for nonlinear Kalman filters

### v10

@@ B.2 Terrain-referenced navigation
This example is adapted from Example 3.3.2 in [26]. As shown in Fig. 11(b), a plane flies above an uneven landscape. The states are the planes’ positions along the two axes, denoted as $x_{1}$ and $x_{2}$. The measurement $y_{k}$ is the terrain’s elevation at the plane’s present location, which, in practice, is calculated as the difference between the barometer and radar altimeter measurements. The plane knows the contour map in advance, so the terrain’s elevation can tell the aircraft its possible locations. However, this task is difficult because all the points on a contour line have the same elevation, and system knowledge must be combined for accurate state estimations. The system inputs are the plane’s speeds along the x and y axes, which are 0.5 km/s and 0, respectively. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (28, 29). Note that all the units for distance in this example are kilometers, and the noises are omitted in the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+0.5\\ x_{2,k-1}\end{bmatrix}$ (28)
$y_{k}=\sin{\sqrt{\big{(}\frac{x_{1,k}}{40}\big{)}^{2}+\big{(}\frac{x_{2,k}}{40}\big{)}^{2}}}$ (29)
@@ B.3 Synchronous generator state estimation
This example is adapted from a research paper [27]. The system model is precisely the same as in the paper. The states of the system are the rotor angle $\delta$, rotor speed $\Delta\omega$, the q-axis component of the voltage $e^{\prime}_{q}$, and the d-axis component of the voltage $e^{\prime}_{d}$. For simplicity, these states are denoted as $x_{1}$ to $x_{4}$. The measurement $y_{k}$ is electrical output power. The system’s inputs include the mechanical input torque $T_{m}$, the steady-state internal voltage of the armature $E_{fd}$, and the terminal bus voltage $V_{t}$. For simplicity, these inputs are denoted as $u_{1}$ to $u_{3}$. The time interval is set to $\Delta t=0.1$ ms. The input profiles are the same as in the paper, which are:
$\begin{cases}u_{1,k}=0.8\\ u_{2,k}=2.11+0.0002k\\ u_{3,k}=1.002\end{cases}$ (30)
All the model parameters are selected to be the same as in [27]. The system’s state transition and measurement functions can be written as (31, 32). Note that the noises are omitted from the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ x_{4,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}\\ x_{2,k-1}\\ x_{3,k-1}\\ x_{4,k-1}\end{bmatrix}$ (31)
$+\begin{bmatrix}377x_{2,k-1}\Delta t\\ \frac{\Delta t}{13}\big{[}u_{1,k-1}-\frac{u_{3,k-1}x_{3,k-1}\sin{x_{1,k-1}}}{0.375}\big{]}\\ \frac{\Delta t}{0.131}[u_{2,k-1}-x_{3,k-1}]\\ \frac{-x_{4,k-1}\Delta t}{0.0131}\end{bmatrix}$
$+\begin{bmatrix}0\\ \frac{\Delta t}{13}\big{[}0.9215u_{3,k-1}^{2}\sin(2x_{1,k-1})-0.05x_{2,k-1}\big{]}\\ \frac{-4.4933\Delta t}{0.131}(x_{3,k-1}-u_{3,k-1}\cos{x_{1,k-1}})\\ \frac{0.6911u_{3,k-1}\sin{x_{1,k-1}}\Delta t}{0.0131}\end{bmatrix}$
$y_{k}=\frac{u_{3,k-1}x_{3,k}\sin{x_{1,k}}}{0.375}+0.9215u_{3,k-1}^{2}\sin(2x_{1,k})$ (32)
The true initial states of the system are $x_{1,0}=0.4,x_{2,0}=x_{3,0}=x_{4,0}=0$. The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $10^{-4}$, $10^{-10}$, $10^{-4}$, and $10^{-4}$. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are $10^{-10}$, $10^{-16}$, $10^{-10}$, and $10^{-10}$.
@@ B.4 Pendulum state estimation
As shown in Fig. 11(c), one end of the pendulum is fixed to the ceiling with a thin rope. The system’s states are the pendulum’s angular position $\theta$ and angular speed $\omega$. The measurement $y_{k}$ is the horizontal component of tension in the rope. The system input is the external torque exerted on the pendulum, which is zero in this example. The mass of the ball is $m=1$ kg, the length of the rope is $l=1$ m, and the gravity acceleration is $g=9.8$m/s2. The time interval is set to $\Delta t=0.01$s. According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (33,34). Note that the noises are omitted in the equations.
$\begin{bmatrix}\omega_{k}\\ \theta_{k}\end{bmatrix}=\begin{bmatrix}\omega_{k-1}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \theta_{k-1}+\omega_{k-1}\Delta t\end{bmatrix}$ (33)
$y_{k}=mg\cos{\theta_{k}}\sin{\theta_{k}}+ml\omega_{k}^{2}\sin{\theta_{k}}$ (34)
@@ B.5 Battery state estimation
This example is adapted from a research paper [28]. As shown in Fig. 11(d), a battery can be modeled as several circuit components, including a voltage source OCV (stands for the cell’s open-circuit voltage), an internal resistor $R_{1}$, a transfer resistor $R_{2}$, and a capacitor $C_{1}$. A battery’s two most important states are its state of charge (SOC) and state of health (SOH). The SOC is the ratio of the battery’s remaining capacity to its present maximum capacity, and the SOH is the ratio of the battery’s present maximum capacity to its initial maximum capacity. In general, SOC describes the battery’s energy level, while SOH describes the battery’s aging level. The states of this system are the SOC, SOH, and capacitor voltage $U_{c}$. The measurement $y_{k}$ is the battery’s terminal voltage. The input is the charging current, which, in this example, is a three-level square wave shown below:
$I_{k}=\begin{cases}-2,75\geq k>15\\ 2,165\geq k>105\\ 0,\text{else}\end{cases}$ (35)
The battery’s initial maximum capacity is $Q_{0}=1$ Ah. The parameters in the equivalent circuit model are chosen to be $R_{1}=0.01,R_{2}=0.05,1/(R_{2}C_{1})=0.008$. The OCV is a function of SOC and SOH, and it can be written as:
$OCV=\sum_{i=0}^{9}\big{(}\frac{SOH-0.8}{0.2}a_{100,i+1}+\frac{1-SOH}{0.2}a_{80,i+1}\big{)}$ (36)
where $a_{100}=[1390.38,-6961.31,14760.31,-17230.92,$ 12055.71,5162.75,1330.60,196.37,15.60,2.96]12055.71,-5162.75,1330.60,-196.37,15.60,2.96]12055.71 , - 5162.75 , 1330.60 , - 196.37 , 15.60 , 2.96 ], $a_{80}=$ [813.94,4229.96,9345.49,11415.38,8396.15,3801.07,[813.94,-4229.96,9345.49,-11415.38,8396.15,-3801.07,[ 813.94 , - 4229.96 , 9345.49 , - 11415.38 , 8396.15 , - 3801.07 , 1043.09,165.29,14.28,2.96]1043.09,-165.29,14.28,2.96]1043.09 , - 165.29 , 14.28 , 2.96 ]. The time interval is set to $\Delta t=1$ s, and the total simulation time is 180 seconds. The system’s state transition and measurement functions can be written as (37,38). Note that the noises are omitted in the equations.
@@
$+\begin{bmatrix}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \big{[}1-\exp{\big{(}\frac{-\Delta t}{R_{2}C_{1}}\big{)}}\big{]}R_{2}I_{k-1}\\ 0\end{bmatrix}$
$y_{k}=OCV(SOC_{k},SOH_{k})+U_{c,k}+R_{1}I_{k}$ (38)
@@ Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
The detailed algorithms of EKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 35. Note that all these algorithms degrade to the conventional framework when the “Recalibrate” and “Back out” steps are replaced by (10).
@@
15:     $\bm{\hat{y}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{i}$
16:     $\bm{P}_{xy,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
17:     $\bm{P}_{y,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
18:      $+\frac{1}{2}\big{[}\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k-1}^{*}\bm{P}_{k|k-1})\big{]}_{ij}$
19:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
20:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
21:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
@@
19:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
20:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k-1}$
21:     $\bm{P}_{xy,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
22:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
23:     $\bm{P}_{y,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}$
24:         $-\bm{\hat{y}}_{k|k-1}]^{T}$
25:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
26:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
27:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$
@@
31:     $\bm{\mathcal{Y}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
32:     $\bm{\hat{y}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Y}}_{i,k|k}$
33:     $\bm{P}_{xy,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
34:     $\bm{P}_{y,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}][\bm{\mathcal{Y}}_{i,k|k}-\bm{\hat{y}}_{k|k}]^{T}$
35:     $\bm{S}_{k|k}=\bm{P}_{y,k|k}+\bm{R}_{k}$
36:     $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}\bm{S}_{k|k}\bm{K}^{T}-\bm{P}_{xy,k|k}\bm{K}^{T}-\bm{K}\bm{P}_{xy,k|k}^{T}$
@@
3:$\bm{\hat{x}}_{0|0}=\operatorname{\mathbb{E}}[\bm{x}_{0}]$
4:$\bm{P}_{0|0}=\operatorname{\mathbb{E}}[(\bm{\hat{x}}_{0|0}-\bm{x}_{0})(\bm{\hat{x}}_{0|0}-\bm{x}_{0})^{T}]$
5:$\lambda=\alpha^{2}(n_{x}+\kappa)-n_{x}$
6:$W_{i}^{(c)}=W_{i}^{(m)}=1/(2n_{x}+2\lambda),i=1,\ldots,n_{x}$
7:$W_{0}^{(c)}=\lambda/(n_{x}+\lambda)+1-\alpha^{2}+\beta$
8:$W_{0}^{(m)}=\lambda/(n_{x}+\lambda)$
@@
11:     $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=chol(\bm{P}_{k-1|k-1})$, where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
12:     $[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=i^{th}\text{ column of }\bm{L},i=1,...,n_{x}$
@@
24:     $\bm{\mathcal{Y}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
25:     $\bm{\hat{y}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Y}}_{i,k|k-1}$
26:     $\bm{P}_{xy,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
27:     $\bm{P}_{y,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}][\bm{\mathcal{Y}}_{i,k|k-1}-\bm{\hat{y}}_{k|k-1}]^{T}$
28:     $\bm{S}_{k|k-1}=\bm{P}_{y,k|k-1}+\bm{R}_{k}$
29:     $\bm{K}=\bm{P}_{xy,k|k-1}\bm{S}_{k|k-1}^{-1}$
30:     $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}(\bm{z}_{k}-\bm{\hat{y}}_{k|k-1})$

### v11

@@ B.2 Terrain-referenced navigation
This example is adapted from Example 3.3.2 in [38]. As shown in Fig. 14(b), a plane flies above an uneven landscape. The states are the planes’ positions along the two axes, denoted as $x_{1}$ and $x_{2}$. The measurement $z_{k}$ is the terrain’s elevation at the plane’s present location, which, in practice, is calculated as the difference between the barometer and radar altimeter measurements. The plane knows the contour map in advance, so the terrain’s elevation can tell the aircraft its possible locations. However, this task is difficult because all the points on a contour line have the same elevation, and system knowledge must be combined for accurate state estimations. The system inputs are the plane’s speeds along the x and y axes, which are 0.5 km/s and 0, respectively. When the time interval is set to $\Delta t=1$s, the system’s state transition and measurement functions can be written as (44) and (45). Note that all the units for distance in this example are kilometers, and the noises are omitted in the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}+0.5\\ x_{2,k-1}\end{bmatrix}$, (44)
$z_{k}=\sin{\sqrt{\big(\frac{x_{1,k}}{40}\big)^{2}+\big(\frac{x_{2,k}}{40}\big)^{2}}}.$ (45)
@@ B.3 Synchronous generator state estimation
This example is adapted from a research paper [39]. The system model is precisely the same as in the paper. The states of the system are the rotor angle $\delta$, rotor speed $\Delta\omega$, the q-axis component of the voltage $e^{\prime}_{q}$, and the d-axis component of the voltage $e^{\prime}_{d}$. For simplicity, these states are denoted as $x_{1}$ to $x_{4}$. The measurement $z_{k}$ is electrical output power. The system’s inputs include the mechanical input torque $T_{m}$, the steady-state internal voltage of the armature $E_{fd}$, and the terminal bus voltage $V_{t}$. For simplicity, these inputs are denoted as $u_{1}$ to $u_{3}$. The time interval is set to $\Delta t=0.1$ ms. The input profiles are the same as in the paper, which are:
$\begin{cases}u_{1,k}=0.8,\\ u_{2,k}=2.11+0.0002k,\\ u_{3,k}=1.002.\end{cases}$ (46)
All the model parameters are selected to be the same as in [39]. The system’s state transition and measurement functions can be written as (47) and (48). Note that the noises are omitted from the equations.
$\begin{bmatrix}x_{1,k}\\ x_{2,k}\\ x_{3,k}\\ x_{4,k}\end{bmatrix}=\begin{bmatrix}x_{1,k-1}\\ x_{2,k-1}\\ x_{3,k-1}\\ x_{4,k-1}\end{bmatrix}+\begin{bmatrix}377x_{2,k-1}\Delta t\\ \frac{\Delta t}{13}\big[u_{1,k-1}-\frac{u_{3,k-1}x_{3,k-1}\sin{x_{1,k-1}}}{0.375}\big]\\ \frac{\Delta t}{0.131}[u_{2,k-1}-x_{3,k-1}]\\ \frac{-x_{4,k-1}\Delta t}{0.0131}\end{bmatrix}$ (47)
$+\begin{bmatrix}0\\ \frac{\Delta t}{13}\big[0.9215u_{3,k-1}^{2}\sin(2x_{1,k-1})-0.05x_{2,k-1}\big]\\ \frac{-4.4933\Delta t}{0.131}(x_{3,k-1}-u_{3,k-1}\cos{x_{1,k-1}})\\ \frac{0.6911u_{3,k-1}\sin{x_{1,k-1}}\Delta t}{0.0131}\end{bmatrix}$,
$z_{k}=\frac{u_{3,k-1}x_{3,k}\sin{x_{1,k}}}{0.375}+0.9215u_{3,k-1}^{2}\sin(2x_{1,k}).$ (48)
The true initial states of the system are $x_{1,0}=0.4,x_{2,0}=x_{3,0}=x_{4,0}=0$. The initial state covariance matrix $P_{0}$ is diagonal, and its diagonal elements are $^{-4}, 1010$$^{-10}$, $10^{-4}$, and $10^{-4}$. The initial state estimations are generated randomly from the multivariate Gaussian distribution with covariance $P_{0}$ and mean equal to the true initial states. The process noise covariance matrix is also diagonal, and its diagonal values are $10^{-10}$, $^{-16}, 1010$$^{-10}$, and $10^{-10}$.
@@ B.4 Pendulum state estimation
As shown in Fig. 14(c), one end of the pendulum is fixed to the ceiling with a thin rope. The system’s states are the pendulum’s angular position $\theta$ and angular speed $\omega$. The measurement $z_{k}$ is the horizontal component of tension in the rope. The system input is the external torque exerted on the pendulum, which is zero in this example. The mass of the ball is $m=1$ kg, the length of the rope is $l=1$ m, and the gravity acceleration is g=9.8g=9.8m/s2. The time interval is set to $\Delta t=0.01$s. According to Newton’s law, when the time interval is small, the system’s state transition and measurement functions can be written as (49) and (50). Note that the noises are omitted in the equations.
$\begin{bmatrix}\omega_{k}\\ \theta_{k}\end{bmatrix}=\begin{bmatrix}\omega_{k-1}-\frac{g}{l}\sin(\theta_{k-1})\Delta t\\ \theta_{k-1}+\omega_{k-1}\Delta t\end{bmatrix}$, (49)
$z_{k}=mg\cos{\theta_{k}}\sin{\theta_{k}}+ml\omega_{k}^{2}\sin{\theta_{k}}.$ (50)
@@ B.5 Battery state estimation
This example is adapted from a research paper [40]. As shown in Fig. 14(d), a battery can be modeled as several circuit components, including a voltage source OCV (stands for the cell’s open-circuit voltage), an internal resistor $R_{1}$, a transfer resistor $R_{2}$, and a capacitor $C_{1}$. A battery’s two most important states are its state of charge (SOC) and state of health (SOH). The SOC is the ratio of the battery’s remaining capacity to its present maximum capacity, and the SOH is the ratio of the battery’s present maximum capacity to its initial maximum capacity. In general, SOC describes the battery’s energy level, while SOH describes the battery’s aging level. The states of this system are the SOC, SOH, and capacitor voltage $U_{c}$. The measurement $z_{k}$ is the battery’s terminal voltage. The input is the charging current, which, in this example, is a three-level square wave shown below:
$I_{k}=\begin{cases}-2,&15<k\leq 75,\\ 2,&105<k\leq 165,\\ 0,&\text{else}.\end{cases}$ (51)
The battery’s initial maximum capacity is $Q_{0}=1\,$Ah. The parameters in the equivalent circuit model are chosen to be $R_{1}=0.01\,\Omega,R_{2}=0.05\,\Omega,1/(R_{2}C_{1})=0.008\,\mathrm{s}$. The OCV is a function of SOC and SOH, and it can be written as:
$OCV=\sum_{i=0}^{9}\big(\frac{SOH-0.8}{0.2}a_{100,10-i}+\frac{1-SOH}{0.2}a_{80,10-i}\big)SOC^{i},$ (52)
where $a_{100}=[1390.38,-6961.31,14760.31,-17230.92,$ 12055.71,5162.75,1330.60,196.37,15.60,2.96]12055.71,-5162.75,1330.60,-196.37,15.60,2.96], $a_{80}=$ [813.94,4229.96,9345.49,11415.38,8396.15,3801.07,[813.94,-4229.96,9345.49,-11415.38,8396.15,-3801.07, 1043.09,165.29,14.28,2.96]1043.09,-165.29,14.28,2.96]. The time interval is set to $\Delta t=1$s, and the total simulation time is 180 seconds. The system’s state transition and measurement functions can be written as (53,54). Note that the noises are omitted in the equations.
@@
$+\begin{bmatrix}I_{k-1}\Delta t\\ \big[1-\exp{\big(\frac{-\Delta t}{R_{2}C_{1}}\big)}\big]R_{2}I_{k-1}\\ 0\end{bmatrix}$
$z_{k}=OCV(SOC_{k},SOH_{k})+U_{c,k}+R_{1}I_{k}$ (54)
@@ Appendix C The detailed algorithm of various nonlinear Kalman filters used in the paper
The detailed algorithms of EKF, IEKF, EKF2, UKF, and CKF with the new framework are shown in Algorithms 26. Note that all these algorithms degrade to the conventional framework when the “Recalibrate” and “Back out” steps are replaced by (13).
@@
2:Initialization:
[…]
5:for every time step $k$ do
6:Predict:
8:  $\bm{\hat{x}}_{k|k-1}=\bm{f}(\bm{\hat{x}}_{k-1|k-1},\bm{u}_{k-1})$
9:  $\bm{P}_{k|k-1}=\bm{F}_{k}\bm{P}_{k-1|k-1}\bm{F}_{k}^{T}+\bm{Q}_{k}$
10:Update:
11:  $\bm{x}_{0}=\bm{\hat{x}}_{k|k-1}$
12:  for $i=1$ to 1000 do
13:   $\bm{H}_{k|k-1}=\frac{\partial\bm{h}}{\partial\bm{x}}|_{\bm{x}_{i-1}}$
14:   $\bm{\hat{z}}_{k|k-1}=\bm{h}(\bm{x}_{i-1})+\bm{H}_{k|k-1}(\bm{x}_{i-1}-\bm{\hat{x}}_{k|k-1})$
15:   $\bm{P}_{xz,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
16:   $\bm{P}_{z,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
17:   $\bm{S}_{k|k-1,i}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
18:   $\bm{K}_{k,i}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1,i}^{-1}$
19:   $\bm{x}_{i}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k,i}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
[…]
21:     $\bm{S}_{k|k-1,i}=\bm{S}_{k|k-1,i-1}$
22:     $\bm{K}_{k,i}=\bm{K}_{k,i-1}$
23:     $\bm{x}_{i}=\bm{x}_{i-1}$
24:     break
25:   end if
26:   if $\max_{j}|1-\bm{x}_{i}(j)/\bm{x}_{i-1}(j)|<0.001$ then
27:     break
28:   end if
29:  end for
30:  $\bm{S}_{k|k-1}=\bm{S}_{k|k-1,i}$
31:  $\bm{K}_{k}=\bm{K}_{i}$
32:  $\bm{\hat{x}}_{k|k}=\bm{x}_{i}$
33:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}-\bm{K}_{k}\bm{S}_{k|k-1}\bm{K}_{k}^{T}$
34:end for
35:Note that rows 16–21 are added to prevent the state estimations from diverging. Namely, when the sign of divergence is detected (row 16), the newest updates of the states, Kalman gain, and measurement residual covariance matrix are withdrawn (rows 17–19). If such a withdrawal is not made, we find in the simulation that the estimation accuracy can be much lower. This characteristic also suggests that the algorithm is not very robust and stable.
@@
15:  $\bm{\hat{z}}_{k|k-1}=\bm{h}(\bm{\hat{x}}_{k|k-1})+\frac{1}{2}\big[\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1})\big]_{i}$
16:  $\bm{P}_{xz,k|k-1}=\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
17:  $\bm{P}_{z,k|k-1}=\bm{H}_{k|k-1}\bm{P}_{k|k-1}\bm{H}_{k|k-1}^{T}$
18:  $\frac{1}{2}\big[\text{tr}(\bm{H}_{i,k|k-1}^{*}\bm{P}_{k|k-1}\bm{H}_{j,k|k-1}^{*}\bm{P}_{k|k-1})\big]_{ij}$
19:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
20:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
21:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
@@
19:  $\bm{\mathcal{Z}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=1,\ldots,2n_{x}$
20:  $\bm{\hat{z}}_{k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Z}}_{i,k|k-1}$
21:  $\bm{P}_{xz,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}$
22:   $\bm{\hat{z}}_{k|k-1}]^{T}$
23:  $\bm{P}_{z,k|k-1}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}$
24:   $\bm{\hat{z}}_{k|k-1}]^{T}$
25:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
26:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
27:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$
@@
31:  $\bm{\mathcal{Z}}_{i,k|k}=\bm{h}(\bm{\mathcal{X}}_{i,k|k}),i=1,\ldots,2n_{x}$
32:  $\bm{\hat{z}}_{k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}\bm{\mathcal{Z}}_{i,k|k}$
33:  $\bm{P}_{xz,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{X}}_{i,k|k}-\bm{\hat{x}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
34:  $\bm{P}_{z,k|k}=\sum_{i=1}^{2n_{x}}\frac{1}{2n_{x}}[\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}][\bm{\mathcal{Z}}_{i,k|k}-\bm{\hat{z}}_{k|k}]^{T}$
35:  $\bm{S}_{k|k}=\bm{P}_{z,k|k}+\bm{R}_{k}$
36:  $\bm{P}_{k|k}=\bm{P}_{k|k-1}+\bm{K}_{k}\bm{S}_{k|k}\bm{K}_{k}^{T}-\bm{P}_{xz,k|k}\bm{K}_{k}^{T}-\bm{K}_{k}\bm{P}_{xz,k|k}^{T}$
@@
[…]
@@
11:  $\bm{L}=[\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=\text{chol}(\bm{P}_{k-1|k-1})$, where $\bm{L}\bm{L}^{T}=\bm{P}_{k-1|k-1}$
12:  $\bm{L}_{1},\ldots,\bm{L}_{n_{x}}]=i^{th}\text{ column of }\bm{L},i=1,...,n_{x}$
@@
24:  $\bm{\mathcal{Z}}_{i,k|k-1}=\bm{h}(\bm{\mathcal{X}}_{i,k|k-1}),i=0,\ldots,2n_{x}$
25:  $\bm{\hat{z}}_{k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(m)}\bm{\mathcal{Z}}_{i,k|k-1}$
26:  $\bm{P}_{xz,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{X}}_{i,k|k-1}-\bm{\hat{x}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}]^{T}$
27:  $\bm{P}_{z,k|k-1}=\sum_{i=0}^{2n_{x}}W_{i}^{(c)}[\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}][\bm{\mathcal{Z}}_{i,k|k-1}-\bm{\hat{z}}_{k|k-1}]^{T}$
28:  $\bm{S}_{k|k-1}=\bm{P}_{z,k|k-1}+\bm{R}_{k}$
29:  $\bm{K}_{k}=\bm{P}_{xz,k|k-1}\bm{S}_{k|k-1}^{-1}$
30:  $\bm{\hat{x}}_{k|k}=\bm{\hat{x}}_{k|k-1}+\bm{K}_{k}(\bm{z}_{k}-\bm{\hat{z}}_{k|k-1})$